While snowboarding, Jonah falls down and his cell phone slides out of his pocket.The cell phone has a mass of 129 g. When it leaves Jonah’s pocket, the cell phone slides on the ice with an initial velocity of 7.2 m/s downhill, sliding directly toward a chairlift pole 22 m away. The icy (negligible friction) hill is sloped at a consistent angle of 19°. What is the cell phone’s speed when it hits the pole? (Neglect frictional forces)
1
Expert's answer
2020-12-17T07:26:20-0500
"0.5mV^2=0.5mv^2+mgh\\V^2=v^2+2gl\sin{19}\\V^2=72^2+2(9.8)(22)\sin{19}\\V=73.0\frac{m}{s}"